# Forecasting age distribution of deaths across countries: Life expectancy and annuity valuation

### v2

@@ 5 Comparisons of life expectancy at birth and annuity price
Once we compute the survival probabilities for ages at and beyond 60, the calculation of annuity prices also depends on the value of the zero-coupon bond price. For an $x$-year-old with a benefit, one monetary amount per year is given by
@@
where $B(0,\tau)$ is the $\tau$-year bond price and $_{\tau}p_{x}$ denotes the survival probability.
Table 3 provides an example of the annuity calculations using the Australian data. We compute the best estimate of the annuity prices for different entry ages and maturities for female and male policyholders residing in Australia. We assume the constant interest rate at the value of $\lambda=4.1\%$ as of 18 May 2025 and hence the zero-coupon bond is given as $B(0,\tau)=\exp^{-\lambda\tau}$.
@@
Figure 9 presents dotcharts for the estimated annuity prices of the 24 countries at entry age 60 with a maturity period of 5 years. With a return of one monetary amount, the estimated annuity price is the highest for Japan, while the two countries with the lowest values are Iceland and Hungary.
@@
In Figure 10, we present dotcharts for the estimated annuity prices of the 24 countries at entry age 70 with a maturity period of 20 years. With a return of one monetary amount, the estimated annuity price is relatively the highest for Japan, while the lowest is for Hungary.
@@
To measure forecast uncertainty, we construct the 50-years-ahead pointwise prediction intervals for the life-table death counts. For each of the 50 years, we search for the optimal tuning parameter that achieves the smallest CPD between the empirical coverage probability and nominal coverage probabilities of 80% and 95%. In Table LABEL:tab:4, we present the 80% and 95% pointwise prediction intervals of annuities for different entry ages and maturities, where age + maturity $\leq 110$.
@@
The forecast uncertainties increase in prediction interval width as maturities increase from $T=5$ to $T=30$ for a given age. They also increase as the initial ages when entering contracts increase from 60 to 105 for a given maturity.
Figures 11a and 11b show the 80% and 95% prediction intervals for annuity prices at entry age 60 with 5-year maturity across 24 countries, while Figures 11c and 11d show the results for entry age 70 with 20-year maturity. With a return of one monetary unit, estimates are shown as open circles, with open intervals indicating prediction intervals. Forecast uncertainty is noticeably higher for Iceland and lower for Hungary. With increasing maturity, prediction intervals widen, reflecting greater long-term uncertainty.
@@ 6 Conclusion
For modeling and forecasting age distribution of death counts, we consider the centered log-ratio and cumulative distribution function transformations. Both transformations allow us to bring a constrained object to an unconstrained space, where the linear techniques can be applied. Such a technique is principal component analysis, which can summarize and model variation in the unconstrained data. Using 24 countries’ data, we evaluate the point forecast accuracy (measured by the Kullback-Leibler and Jensen-Shannon divergences) and the interval forecast accuracy (measured by the coverage probability difference between empirical and nominal coverage probabilities and mean interval score). For producing both point and interval forecasts, the cumulative distribution function transformation provides smaller point and interval forecast errors than the centered log-ratio transformation. Between the two ways of selecting the number of principal components, we advocate the selection of $K=6$ for forecasting.
@@
We apply the cumulative distribution function transformation to generate 50-years-ahead forecasts of age-specific life-table death counts for 24 countries. From the forecasted age distribution of death counts, we evaluate and compare the life expectancy at birth across multiple countries. From the forecasted life-table death counts, we convert them into age-specific survival probabilities and obtain temporary annuity prices. As expected, we observe that the survival probability has a pronounced impact on annuity prices. Coupling with the uncertainty in interest rates, we quantify the uncertainty of annuity prices via pointwise prediction intervals. For reproducibility, the computer [Uncaptioned image] code is available at https://github.com/hanshang/CLR_vs_CDF_transformation.
@@
There are a few ways in which the methodology presented in the paper could be extended, and we briefly discuss four. 1) When forecasting principal component scores, we implement exponential smoothing. Other univariate or multivariate time-series forecasting methods may also be plausible. 2) We consider two transformations, namely the centered log-ratio and cumulative distribution function transformations. There exist other transformations that could also be considered, namely additive or isometric log-ratio (Greenacre et al., 2023). While these transformations are variants of the centered log-ratio, we could also consider intrinsic statistical models, such as Wasserstein autoregression (Zhang et al., 2022) or optimal transport (Zhu & Müller, 2023), for modeling density-valued time series. 3) Although we calculate temporary annuities, other types of annuity prices, such as the whole-life or deferred annuity, are possible. 4) We could consider cohort life-table death counts for modeling a particular group of individuals, such as baby boomers.
@@
The primary source for the analysis is the freely available data in the Human Mortality Database (HMD) that is publicly accessible subject to creating a free HMD account. No further pre-processing was applied to this data. For reproducibility, the computer [Uncaptioned image] code is available at https://github.com/hanshang/CLR_vs_CDF_transformation.
@@
This research was supported by the Australian Research Council Future Fellowship (grant no. FT240100338).

### v3

@@ 6.2 Comparison of annuity price
Once we compute the survival probabilities for ages 60 and beyond, the calculation of annuity prices also depends on the zero-coupon bond price. For an $x$-year-old with a benefit, a monetary amount per year is given by
@@
where $B(0,\tau)$ is the price of the $\tau$-year bond and $_{\tau}p_{x}$ denotes the survival probability.
Table 4 provides an example of annuity calculations using Australian data. We calculate the optimal estimate of annuity prices for various entry ages and maturities for female and male policyholders residing in Australia. We assume a constant interest rate at the value of $\lambda=4.1\%$ as of 18 May 2025, and hence the zero-coupon bond is given as $B(0,\tau)=\exp^{-\lambda\tau}$.
@@
Figure 13 shows dot charts for the estimated annuity prices of the 24 countries at entry age 60 with a maturity period of 5 years. With a return of one monetary amount, the estimated annuity price is the highest for Japan, while the two countries with the lowest values are Iceland and Hungary.
@@
In Figure 14, we present dot charts of the estimated annuity prices for the 24 countries at entry age 70 with a 20-year maturity. With a return of one monetary amount, the estimated annuity price is highest in Japan and lowest in Hungary.
@@
To measure forecast uncertainty, we construct the 50-years-ahead pointwise prediction intervals for the life-table death counts. For each of the 50 years, we search for the optimal tuning parameter that minimizes the CPD between the empirical coverage probability and the nominal coverage probabilities of 80% and 95%. In Table LABEL:tab:4, we present the 80% and 95% pointwise prediction intervals of annuities for different entry ages and maturities, where age + maturity $\leq 110$.
@@
The forecast uncertainty increases the width of the prediction interval as maturity increases from $T=5$ to $T=30$ for a given age. They also increase as the initial ages at contract entry range from 60 to 105 for a given maturity. Figures 15a and 15b show the 80% and 95% prediction intervals for annuity prices at entry age 60 with 5-year maturity in 24 countries, while Figures 15c and 15d show the results for entry age 70 with 20-year maturity.
@@
With a return of one monetary unit, estimates are shown as open circles, with open intervals indicating 80% and 95% pointwise prediction intervals in red and blue, respectively. The forecast uncertainty is noticeably higher for Iceland and lower for Hungary. As maturity increases, the prediction intervals widen, reflecting greater long-term uncertainty.
@@ 7 Conclusion
We present a comprehensive cross-country validation study of compositional mortality forecasting methods. For modeling and forecasting the age distribution of death counts, we consider the centered log-ratio and cumulative distribution function transformations. Both transformations allow us to map a constrained object into an unconstrained space, where linear techniques can be applied. Such a technique is principal component analysis, which can summarize and model variation in the unconstrained data. Between the two transformations, we demonstrate an additional use of the CDF transformation, providing a scale-free measure to compare the probabilities of dying across genders and countries.
Using data from 24 countries, we evaluate point forecast accuracy (measured by the Kullback-Leibler and Jensen-Shannon divergences) and interval forecast accuracy (measured by the coverage probability difference between empirical and nominal coverage probabilities and the mean interval score). For producing both point and interval forecasts, the cumulative distribution function transformation systematically provides smaller point and interval forecast errors than the centered log-ratio transformation. Between the two methods for selecting the number of principal components, we advocate $K=6$ for forecasting.
@@
We apply the cumulative distribution function transformation to generate 50-years-ahead forecasts of age-specific life-table death counts for 24 countries. From the forecasted age distribution of death counts, we evaluate and compare the life expectancy at birth across multiple countries. From the forecasted life-table death counts, we convert them into age-specific survival probabilities and obtain temporary annuity prices. As expected, we observe that the survival probability has a pronounced impact on actuarially fair annuity prices, reflecting gender gap and cross-country heterogeneity. Coupling with the current interest rates in Table 1, we quantify the uncertainty of annuity prices via pointwise prediction intervals. These prediction intervals primarily reflect model uncertainty rather than uncertainty around interest rates themselves. For reproducibility, the computer [Uncaptioned image] code is available at https://github.com/hanshang/CLR_vs_CDF_transformation.
@@
There are several ways the methodology presented in the paper could be extended, and we briefly discuss six. 1) We consider a univariate functional time-series method to forecast each series individually. While it provides a common approach for comparing the two transformations, other joint modeling methods could also be used. Also, when forecasting principal component scores, we implement exponential smoothing. Other time-series forecasting methods may also be used. 2) We consider two transformations, namely the centered log-ratio and cumulative distribution function transformations. There exist other transformations that could be considered, namely additive or isometric log-ratio (Greenacre et al., 2023). While these transformations are variants of the centered log-ratio, we could also consider intrinsic statistical models, such as Wasserstein autoregression (Zhang et al., 2022) or optimal transport (Zhu and Müller, 2023), for modeling density-valued time series. 3) We could also consider testing for time-stochastic dominance between populations, as studied by Lee et al. (2023). 4) For all countries, the sample period encompasses the COVID-19 pandemic, but we do not examine the changes in the age distribution of deaths before and after the pandemic. 5) Although we calculate temporary annuities, other types of annuity prices, such as the whole-life or deferred annuity, are possible. 6) We could consider cohort life-table death counts for modeling a particular group of individuals, such as baby boomers.
@@
The primary source for the analysis is the freely available data in the Human Mortality Database (HMD), which is publicly accessible after creating a free HMD account.
@@
This research was supported by the Australian Research Council Future Fellowship (grant no. FT240100338). The first author is grateful for the insightful comments and suggestions from the seminar participants at the University of New South Wales.